jamie1837:

Helena uses a mixture of compost and topsoil in her garden. She purchased a total of 10 cubic yards of compost and topsoil for $180. If compost costs $25 per cubic yard and topsoil costs $15 per cubic yard, how many cubic yards of topsoil did she purchase?

Astro:

First: We know that each compost costs $25=25C Topsoil costs $15=15T Then we know the price of compost plus the cost of topsoil=10. What do you think we would do next?

Astro:

We also know C+T=10 25C+15T=180

Astro:

So for C it would be C=10-T

Astro:

Does that make sense @jamie1837

Sock:

Also welcome to Qc jamie!